# Dissipative dynamics of an open quantum battery in the BTZ spacetime

### v1

@@ 4 Discussions and Conclusions
In summary, we studied how charging performances of a quantum battery, modeled as a two-level system, are influenced by the presence of vacuum fluctuations of a quantum field satisfying the Dirichlet, transparent, and Neumann boundary conditions in the BTZ spacetime. We obtained the analytical dynamics of this QB, and analytically calculated the average energy stored in the QB. We find that when the driving amplitude is stronger/weaker than the energy-level spacing of the QB the pure dephasing dissipative coupling results in better/worse charging performance than the decoherence dissipative coupling case. We also found that compared with the closed QB case, as a result of dissipative coupling the higher Hawking temperature helps to improve the charging performance under certain conditions. This means that it is possible for us to extract energy from vacuum fluctuations in curved spacetime via dissipation in charging protocol. We also considered how different boundary conditions for quantum field affect the charging performance. We found that the boundary condition not only quantitatively but also qualitatively affects the the QB dynamics. It is shown that the maximum average energy stored in the QB increases with the decrease of Hawking temperature for the Dirichlet boundary case, while it is opposite for the transparent and Neumann boundary cases. Furthermore, we also addressed the charging stability by monitoring the energy behaviour after the charging protocol has been switched off. We found that from the perspective of the charging stability, the pure dephasing dissipative coupling may lead to better/worse charging performance than the decoherence dissipative coupling case when the driving amplitude is stronger/weaker than the energy-level spacing of the QB. This result is valid for all the three boundary condition cases. Our study presents a general framework to investigate relaxation effects in curved spacetime, and reveals how spacetime properties and filed boundary condition affect the charging process. This in turn, from a simple quantum device for the production, storage, and transfer of energy, may shed light on the exploration of the spacetime properties and thermodynamics via the charging protocol.

### v4

@@ 3.5 Charging stability
4 Quantum battery in AdS spacetime
Since the BTZ black hole spacetime can be obtained from identifications of $(2+1)$-dimensional AdS spacetime, so does the relevant Wightman function of quantum field shown above, it is interesting to see whether there is difference of the charging performance for these two spacetimes. In our case, $(2+1)$-dimensional AdS with cosmological constant $\Lambda=-1/\ell^{2}$ can be expressed as the hyperboloid $X^{2}_{1}+X^{2}_{2}-T^{2}_{1}-T^{2}_{2}=-\ell^{2}$, embedded in the $(2+2)$-dimensional flat spacetime $dS^{2}=dX^{2}_{1}+dX^{2}_{2}-dT^{2}_{1}-dT^{2}_{2}$, with $\ell$ being the AdS-length. We will consider the AdS-Rinder metric
$ds^{2}=-\bigg{(}\frac{r^{2}}{\ell^{2}}-1\bigg{)}dt^{2}+\bigg{(}\frac{r^{2}}{\ell^{2}}-1\bigg{)}^{-1}dr^{2}+r^{2}d\phi^{2},$ (47)
which is obtained via the transformations:
$T_{1}=\ell\sqrt{\frac{r^{2}}{\ell^{2}}}\cosh\phi,~{}~{}X_{1}=\ell\sqrt{\frac{r^{2}}{\ell^{2}}}\sinh\phi,~{}~{}T_{2}=\ell\sqrt{\frac{r^{2}}{\ell^{2}}-1}\sinh\frac{t}{\ell},~{}~{}X_{2}=\ell\sqrt{\frac{r^{2}}{\ell^{2}}-1}\cosh\frac{t}{\ell}.$
Note that an accelerate horizon presents at $r_{\text{h}}=\ell$. Correspondingly, the vacuum Wightman function for a (massless) conformally coupled scalar field in $(2+1)$-dimensional AdS reads
$G^{+}_{\text{AdS}}(x,x^{\prime})=\frac{1}{4\pi\sqrt{2}\ell}\bigg{(}\frac{1}{\sqrt{\sigma(x,x^{\prime})}}-\frac{\zeta}{\sqrt{\sigma(x,x^{\prime})+2}}\bigg{)},$ (49)
where $\sigma(x,x^{\prime})=\frac{1}{2\ell^{2}}\big{[}(X_{1}-X^{\prime}_{1})^{2}-(T_{1}-T^{\prime}_{1})^{2}+(X_{2}-X^{\prime}_{2})^{2}-(T_{2}-T^{\prime}_{2})^{2}\big{]}$ denotes the square distance between $x$ and $x^{\prime}$ in the embedding space. Similarly, $\zeta\in\{1,0,-1\}$ respectively specifies the the Dirichlet $(\zeta=1)$, transparent $(\zeta=0)$, and Neumann $(\zeta=-1)$ boundary conditions satisfied by the field at spatial infinity PhysRevD.18.3565 .
Similar to the BTZ black hole case, we also assume that the QB is spatially fixed at a constant $r$ in the AdS spacetime such that $\delta\phi=0$, $\Delta\tau=\tau-\tau^{\prime}=\sqrt{-g_{00}}\Delta\tau$. In this case, inserting the AdS Wightman function from (49) into the QB state parameter (2.1), we can find
$\gamma_{+}$ $=$ $\frac{\mu^{2}\cos^{2}(\theta-\Theta)}{2}\bigg{(}\frac{1}{e^{\Omega/T}+1}\bigg{)}\bigg{[}1-\zeta P_{-\frac{1}{2}+i\frac{\Omega}{2\pi T}}(1+8\pi^{2}\ell^{2}T^{2})\bigg{]},$
$\gamma_{-}$ $=$ $\frac{\mu^{2}\cos^{2}(\theta-\Theta)}{2}\bigg{(}\frac{1}{e^{-\Omega/T}+1}\bigg{)}\bigg{[}1-\zeta P_{-\frac{1}{2}-i\frac{\Omega}{2\pi T}}(1+8\pi^{2}\ell^{2}T^{2})\bigg{]},$
$\gamma_{z}$ $=$ $\frac{\mu^{2}\sin^{2}(\theta-\Theta)}{4}\bigg{[}1-\zeta P_{-\frac{1}{2}}(1+8\pi^{2}\ell^{2}T^{2})\bigg{]},$ (50)
where $P_{\nu}(x)$ denotes the associated Legendre function of the first kind. Therefore, one can find the ratio between the excitation rate and the deexcitation rate is $\gamma_{+}/\gamma_{-}=e^{-\Omega/T}$. It means the QB will feel as if it were immersed into a thermal bath with the temperature $T=r_{\text{h}}/(2\pi\ell\sqrt{r^{2}-r^{2}_{\text{h}}})$, that is defined as the local KMS temperature of the field in AdS spacetime. Note that this temperature $T=r_{\text{h}}/(2\pi\ell\sqrt{r^{2}-r^{2}_{\text{h}}})$ actually is the temperature felt by an accelerated observer in $\mathrm{AdS}_{3}$ spacetime and the magnitude of the acceleration is given by $|a|=\frac{r}{\ell\sqrt{r^{2}-\ell^{2}}}$ HENDERSON2020135732 , where $r$ denotes the location of the observer. The minimum acceleration is $|a|=1/\ell$ which happens near $r=\infty$, while $|a|\rightarrow\infty$ when $r\rightarrow\ell$. It is demonstrated in Ref. Jennings_2010 that there exists a critical acceleration, $a_{c}=1/\ell$, in $\mathrm{AdS}_{3}$ spacetime, and only the observer with a constant super-critical acceleration $a$ (i.e., $a>1/\ell$) can register the quasi-thermal response with a temperature equal to $T=\sqrt{a^{2}-\ell^{-2}}/(2\pi)$. We will consider this kind of acceleration hereafter.
Repeat the same calculations as that for the BTZ black hole case shown above, we can also find that for the case of decoherence coupling ($\theta=0$ in Eq. (4)) the corresponding average energy associated to the QB in the AdS spacetime is
$\langle E_{\text{QB}}(\tau)\rangle=\frac{\Delta}{2}\bigg{\{}1-\frac{\Delta^{2}}{\Omega^{2}}e^{-\Gamma^{\text{r}}_{0}\tau}-\frac{\Delta}{\Omega}\bigg{(}\frac{e^{\Omega/T}-1}{e^{\Omega/T}+1}\bigg{)}(1-e^{-\Gamma^{\text{r}}_{0}\tau})-\frac{A^{2}}{\Omega^{2}}\cos(\Omega\tau)e^{-\Gamma_{0}\tau}\bigg{\}},$ (51)
where
$\Gamma^{\text{r}}_{0}$ $=$ $\frac{\mu^{2}\Delta^{2}}{2\Omega^{2}}\big{[}1-\zeta P_{-\frac{1}{2}+i\frac{\Omega}{2\pi T}}(1+8\pi^{2}\ell^{2}T^{2})\big{]},$
$\Gamma_{0}$ $=$ $\frac{1}{2}\Gamma^{\text{r}}_{0}+\frac{\mu^{2}A^{2}}{2\Omega^{2}}\big{[}1-\zeta P_{-\frac{1}{2}}(1+8\pi^{2}\ell^{2}T^{2})\big{]}.$ (52)
While for the pure dephasing coupling case ($\theta=\pi/2$ in Eq. (4)), the corresponding average energy associated to the QB is found to be
$\langle E_{\text{QB}}(\tau)\rangle=\frac{\Delta}{2}\bigg{\{}1-\frac{\Delta^{2}}{\Omega^{2}}e^{-\Gamma^{\text{r}}_{\pi/2}\tau}-\frac{\Delta}{\Omega}\bigg{(}\frac{e^{\Omega/T}-1}{e^{\Omega/T}+1}\bigg{)}(1-e^{-\Gamma^{\text{r}}_{\pi/2}\tau})-\frac{A^{2}}{\Omega^{2}}\cos(\Omega\tau)e^{-\Gamma_{\pi/2}\tau}\bigg{\}},$ (53)
where
$\Gamma^{\text{r}}_{\pi/2}$ $=$ $\frac{\mu^{2}A^{2}}{2\Omega^{2}}\big{[}1-\zeta P_{-\frac{1}{2}+i\frac{\Omega}{2\pi T}}(1+8\pi^{2}\ell^{2}T^{2})\big{]},$
$\Gamma_{\pi/2}$ $=$ $\frac{1}{2}\Gamma^{\text{r}}_{\pi/2}+\frac{\mu^{2}\Delta^{2}}{2\Omega^{2}}\big{[}1-\zeta P_{-\frac{1}{2}}(1+8\pi^{2}\ell^{2}T^{2})\big{]}.$ (54)
Comparing the average energies in Eqs. (51) and (53) (for AdS spacetime case) with that in Eqs. (3.1) and (3.1) (for the BTZ black hole case), we can find that their expressions are exactly the same, while only with different incoherent relaxation rate $\Gamma^{\text{r}}_{\theta}$ and the dephasing rate $\Gamma_{\theta}$. Specifically, if we assume the temperature felt by the QB in the BTZ spacetime and AdS spacetime is the same, then we can find the relation of the incoherent relaxation rate (the dephasing rate) in these two different spacetimes is given by
$\Gamma^{\text{r}}_{0\text{(BTZ)}}$ $=$ $\Gamma^{\text{r}}_{0\text{(AdS)}}+\frac{\mu^{2}\Delta^{2}}{\Omega^{2}}\sum^{\infty}_{n=1}\bigg{[}P_{-\frac{1}{2}+i\frac{\Omega}{2\pi T}}(\cosh\alpha_{n})-\zeta P_{-\frac{1}{2}+i\frac{\Omega}{2\pi T}}(\cosh\beta_{n})\bigg{]},$
$\Gamma_{0\text{(BTZ)}}$ $=$ $\Gamma_{0\text{(AdS)}}+\frac{\mu^{2}\Delta^{2}}{2\Omega^{2}}\sum^{\infty}_{n=1}\bigg{[}P_{-\frac{1}{2}+i\frac{\Omega}{2\pi T}}(\cosh\alpha_{n})-\zeta P_{-\frac{1}{2}+i\frac{\Omega}{2\pi T}}(\cosh\beta_{n})\bigg{]}$ (55)
$+\frac{\mu^{2}A^{2}}{\Omega^{2}}\sum^{\infty}_{n=1}\bigg{[}P_{-\frac{1}{2}(\cosh\alpha_{n})}-\zeta P_{-\frac{1}{2}}(\cosh\beta_{n})\bigg{]},$
and
$\Gamma^{\text{r}}_{\pi/2\text{(BTZ)}}$ $=$ $\Gamma^{\text{r}}_{\pi/2\text{(AdS)}}+\frac{\mu^{2}A^{2}}{\Omega^{2}}\sum^{\infty}_{n=1}\bigg{[}P_{-\frac{1}{2}+i\frac{\Omega}{2\pi T}}(\cosh\alpha_{n})-\zeta P_{-\frac{1}{2}+i\frac{\Omega}{2\pi T}}(\cosh\beta_{n})\bigg{]},$
$\Gamma_{\pi/2\text{(BTZ)}}$ $=$ $\Gamma_{\pi/2\text{(AdS)}}+\frac{\mu^{2}A^{2}}{2\Omega^{2}}\sum^{\infty}_{n=1}\bigg{[}P_{-\frac{1}{2}+i\frac{\Omega}{2\pi T}}(\cosh\alpha_{n})-\zeta P_{-\frac{1}{2}+i\frac{\Omega}{2\pi T}}(\cosh\beta_{n})\bigg{]}$ (56)
$+\frac{\mu^{2}\Delta^{2}}{\Omega^{2}}\sum^{\infty}_{n=1}\bigg{[}P_{-\frac{1}{2}(\cosh\alpha_{n})}-\zeta P_{-\frac{1}{2}}(\cosh\beta_{n})\bigg{]}.$
Here the subscripts “BTZ” and “AdS” respectively denote the rates for the BTZ black hole case and AdS spacetime case. Therefore, we can conclude that although the average energies associated to the QB in both the BTZ black hole and AdS spacetime share the same expression, the corresponding incoherent relaxation rate and dephasing rate are different. This means that the average energies associated to the QB in both the BTZ black hole and AdS spacetime share the same behavior with respect to the evolution time, however, their damping strengths are different due to different dissipation from vacuum field fluctuations in different spacetime. Indeed, as shown in (4) and (4), the incoherent relaxation rate (and the dephasing rate) for the AdS spacetime case arises as the $n=0$ term of that for the BTZ black hole case, while the remaining terms ranging from $n=1,\cdots,\infty$ constitute the novel black hole effects. We note that these additional terms are parameterized by the black hole mass, $M$, and has been shown to be with more pronounced effects arising for smaller $M$, such as Fisher information FIBH .
In the infinite time limit, i.e., $\tau\rightarrow\infty$, for both the decoherence and pure dephasing coupling cases in AdS spacetime the average energies in (51) and (53) approach to
$\langle E_{\text{QB}}(\infty)\rangle=\frac{\Delta}{2}\bigg{\{}1-\frac{\Delta}{\Omega}\bigg{(}\frac{e^{\Omega/T}-1}{e^{\Omega/T}+1}\bigg{)}\bigg{\}}.$ (57)
It is shown that in the infinite evolution time limit the average energies of the QB are independent of the kinds of dissipation, while depend on the local KMS temperature of the field in the AdS spacetime. Besides, we can find these average energies are exactly the same as that for the BTZ black hole case shown in (3.1). Therefore, one can not tell the difference between the BTZ black hole and the AdS spacetime in terms of the average energy associated to the QB in the infinite evolution time limit.
@@ 5 Discussions and Conclusions
In summary, we studied how charging performances of a quantum battery, modeled as a two-level system, are influenced by the presence of vacuum fluctuations of a quantum field satisfying the Dirichlet, transparent, and Neumann boundary conditions in the BTZ spacetime. We obtained the analytical dynamics of this QB, and analytically calculated the average energy stored in the QB. We find that when the driving amplitude is stronger/weaker than the energy-level spacing of the QB the pure dephasing dissipative coupling results in better/worse charging performance than the decoherence dissipative coupling case. We also found that compared with the closed QB case, as a result of dissipative coupling the higher local Hawking temperature helps to improve the charging performance under certain conditions. This means that it is possible for us to extract energy from vacuum fluctuations in curved spacetime via dissipation in charging protocol. We also considered how different boundary conditions for quantum field affect the charging performance. We found that the boundary condition not only quantitatively but also qualitatively affects the the QB dynamics. It is shown that the maximum average energy stored in the QB increases with the decrease of the local Hawking temperature for the Dirichlet boundary case, while it is opposite for the transparent and Neumann boundary cases. Furthermore, we also addressed the charging stability by monitoring the energy behaviour after the charging protocol has been switched off. We found that from the perspective of the charging stability, the pure dephasing dissipative coupling may lead to better/worse charging performance than the decoherence dissipative coupling case when the driving amplitude is stronger/weaker than the energy-level spacing of the QB. This result is valid for all the three boundary condition cases. Our study presents a general framework to investigate relaxation effects in curved spacetime, and reveals how spacetime properties and filed boundary condition affect the charging process. This in turn, from a simple quantum device for the production, storage, and transfer of energy, may shed light on the exploration of the spacetime properties and thermodynamics via the charging protocol.
In our model, the two-level system which is coupled a massless scalar field in the BTZ spacetime can be considered as a detector. From the dynamics of this two-level system, we can read out the information of quantum field in the BTZ spacetime, e.g., the field correlation function PhysRevD.98.105011 , and the spacetime curvature PhysRevD.105.125011 . According to the AdS/CFT correspondence theory, actually one can find the corresponding operators in the boundary CFT PhysRevD.73.086003 ; hamilton2007local . In this sense, through appropriate operation on the two-level system, we can obtain the spacetime properties and quantum field properties in BTZ spacetime from the two-level system’s observable, thus obtain the spacetime thermodynamics, singularity, and so on. When applying the AdS/CFT correspondence theory, we can also reveal the corresponding physics in terms of a dual field theory.
Note that our QB model is general, which can be extended to different scenarios where the QB moves along different trajectories in various spacetimes. For example, we can consider the relevant physics of QB in quantum BTZ black hole QBTZ ; universe10090358 , that incorporates the exact backreaction from the fields and whose thermodynamics has been explored recently PhysRevLett.130.161501 ; PhysRevD.109.124040 ; PhysRevD.110.106004 ; mansoori2024 ; PhysRevD.110.024054 ; PhysRevLett.133.181501 . What we need to do in the extension is to replace the Wightman function of the field shown in (23) with the one in the quantum BTZ black hole. Then we repeat the relevant calculations and could analyze how the parameters (or properties) of quantum BTZ black hole affect the charging performances of the QB, as done for that in BTZ black hole. Which in turn could also shed light on new properties of quantum BTZ black hole when comparing the relevant results with that for BTZ black hole case. When considering the backreaction into the spacetime geometry that from the quantum field, there is reason to expect that some new phenomena might emerge in the charging protocol QBTZ ; universe10090358 ; PhysRevLett.130.161501 ; PhysRevD.109.124040 ; PhysRevD.110.106004 ; mansoori2024 ; PhysRevD.110.024054 ; PhysRevLett.133.181501 . Which are left for the future investigation.